OpenStudy (anonymous):

Consider the following point and line. Point Line (4/5, 5/6) 3x + 4y = 0 (a) Write the equation of the line through the point parallel to the given line.

OpenStudy (anonymous):

the slope is -3/4 so (y-5/6) = -3(x-4/5)/4

OpenStudy (anonymous):

What does it mean for two lines to be parallel?

OpenStudy (anonymous):

they never touch? i put it into standard form and got y=-(3/40x

OpenStudy (anonymous):

i mean -3/4

OpenStudy (anonymous):

They never touch, that's true. But they also have the same slope.

OpenStudy (anonymous):

do u know the point slope form of a line tiffany?

OpenStudy (anonymous):

nope?

OpenStudy (anonymous):

\[y-k = m(x-h)\] Where m is the slope and (h,k) is any point on the line.

OpenStudy (anonymous):

So since you know that the line is parallel it will have the same slope of -3/4

OpenStudy (anonymous):

so how do i find out the B which is the x-inter. right?

OpenStudy (anonymous):

y - intercept yeah

OpenStudy (anonymous):

Ok so you have a point on the line you are looking for, and you have the slope (because it's parallel). Plug it into the point-slope formula.

OpenStudy (anonymous):

Then we can solve for the 'slope intercept' form which will give us b.

OpenStudy (anonymous):

Though really the point slope equation is the equation for the line.

OpenStudy (anonymous):

so it should look like y-(5/6)=-3/4(x-5/6)

OpenStudy (anonymous):

Close. Your h isn't quite right.

OpenStudy (anonymous):

y-(5/6) = (-3/4)(x-(4/5))

OpenStudy (anonymous):

You just had 5/6 again instead of the x coordinate of the point.

OpenStudy (anonymous):

ugh this is really confusing..

OpenStudy (anonymous):

No you're doing it right.

OpenStudy (anonymous):

\[y - \frac{5}{6} =- \frac{3}{4}(x-\frac{4}{5})\]

OpenStudy (anonymous):

then would i distribute the -3/4?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

do i subtract -5/6 from both sides.?

OpenStudy (anonymous):

well add

OpenStudy (anonymous):

Yes add 5/6 to both sides

OpenStudy (anonymous):

After you distribute.

OpenStudy (anonymous):

thanks!!!